OpenStudy (anonymous):

will give a metal and become fan to best answer find the LCM. 10x and 35x^5 A. 5x B. 350x^5 C. 70x^5 D. 70x^6

jigglypuff314 (jigglypuff314):

LCM: Lowest Common Multiple basically, what factors do the two values have in common? if you break down each 10x = 2 times 5 times x 35x^5 = 7 times 5 times x times x times x times x times x ^ ^ so they both have a 5 and a x in common right?

OpenStudy (anonymous):

umm...

jigglypuff314 (jigglypuff314):

wait wait wait sorry >,< I had a brain fart gimme a sec to re-explain

OpenStudy (anonymous):

lol, ok :D

jigglypuff314 (jigglypuff314):

LCM: Lowest Common Multiple multiples are like (for example) for 3, multiples are 3, 6, 9, 12, 15, etc Consider LCM of 12 and 20: 12: 2 * 2 * 3 20: 2 * 2 * 5 --------------------- .....2 * 2 * 3 * 5 = 60 = LCM <- shared factors aren't repeated so for things with variables... 10x: 5 * 2 * x 35x^5: 5 * 7 * x * x * x * x * x since they share 5 and x then LCM = 5 * x * 2 * 7 * x * x * x * x =

OpenStudy (anonymous):

ok, so C?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

ok! thanks!